# Learning to Play Pursuit-Evasion with Dynamic and Sensor Constraints

### v1

@@ Abstract
We present a multi-agent reinforcement learning approach to solve a pursuit-evasion game between two players with car-like dynamics and sensing limitations. We develop a curriculum for an existing multi-agent deterministic policy gradient algorithm to simultaneously obtain strategies for both players, and deploy the learned strategies on real robots moving as fast as 2 m/s in indoor environments. Through experiments we show that the learned strategies improve over existing baselines by up to 30% in terms of capture rate for the pursuer. The learned evader model has up to 5% better escape rate over the baselines even against our competitive pursuer model. We also present experiment results which show how the pursuit-evasion game and its results evolve as the player dynamics and sensor constraints are varied. Finally, we deploy learned policies on physical robots for a game between the F1TENTH and JetRacer platforms and show that the learned strategies can be executed on real-robots. Our code and supplementary material including videos from experiments are available at https://gonultasbu.github.io/pursuit-evasion/.
@@ I Introduction
As an example, in the classical homicidal chauffeur game [3], the pursuer is a car-like vehicle whose goal is to hit a pedestrian who is not subject to any turning constraints. In this game, the pursuer has three state variables (position and orientation) and the evader has two (only position). One might view the solution of this game as a subspace (manifold) of this five dimensional joint state space. The tangent space of this solution manifold corresponds to the optimal control actions of the players. In some cases, the differential equations for the evolution of the game when the players play optimally can be derived in closed form. The strategies for a given instance is usually obtained by numerical integration. Alternatively, the state space can be discretized and the solution can be obtained using dynamic programming techniques.
Figure 1: Real-world experiments: Execution of the learned pursuit-evasion policies on two robotic cars with varying wheelbase, velocity and steering angle constraints.
However, such methods become quickly intractable as the dimensionality of the game increases. For example, in his book, Isaacs presents only a partial solution for the “game of two cars” where both players have car-like dynamics. The dimensionality issue becomes more severe when sensing is involved: in the standard formulations, it is assumed that the players can observe each others’ state fully at each time step. In practice, this may not be always true due to sensor and observability limitations. The resulting “partial observability games” have even higher (sometimes infinite) dimensionality as the players now need to reason about all feasible states that are consistent with their observation histories.
Reinforcement Learning (RL) [4] methods provide an alternative approach to construct the solution manifold by learning it through game play. Modern RL methods often represent the solution as a neural network which takes the observations as input and returns a distribution over the actions. In this paper, we investigate whether recently developed multi-agent reinforcement learning (MARL) techniques can solve a challenging game where the pursuer is a car with five state variables including velocities. We investigate two versions of the evader. It is either a car with the same kinematic model or a point mass. Hence, the joint-state dimensionality can be as high as 10. The players are subject to field-of-view sensing restrictions for both the viewing angle and viewing distance (Section III). We cast the game as a competitive Partially Observable Stochastic Game (POSG) and use the MARL algorithm multi-agent deep deterministic policy gradient (MADDPG) [5] to simultaneously train pursuer and evader policies. Our problem formulation and proposed approach do not require any pretrained or designed expert evader policy and generalizes over varying agent constraints.
@@
  • We develop a curriculum strategy for MADDPG to solve a pursuit-evasion game with complex dynamics and sensing constraints. We show that our method is able to train competitive pursuer and evader policies without the need for expert policies.
  • We provide insights on how the pursuit-evasion game and its results evolve as the player dynamic and sensor constraints are varied.
  • In simulation experiments we individually compare pursuer and evader policies against existing baselines.
  • We deploy learned policies on physical robots for a game between the F1TENTH[6] and JetRacer [7] platforms moving at high-speed and showcase a successful policy transfer to real world.
Overall, our results show the potential of multi-agent reinforcement learning in navigating complex pursuit-evasion scenarios within dynamic environments and demonstrate that simulation-trained policies are directly transferable to real autonomous agents.
@@ II-A Pursuit-Evasion
The mathematical origins of the pursuit-evasion problem can be traced back to 1930’s as the Lion and Man problem. In the original formulation, a lion and a human are in a circular area where the lion tries to catch the human as quickly as possible when the human tries to avoid the lion for as long as possible. Since then, various variants of this game have been studied in the robotics literature [8]. In his seminal work [3] Isaacs presented a differential game based formulation to the pursuit-evasion problem which has served as the basis of game-theoretic approaches.
@@
In control-theoretic formulations, both the pursuer and evader in a pursuit-evasion problem have constraints on maximum velocities. One prominent example is the Homicidal chauffeur problem formulated by Isaacs [3], which is a special case of the pursuit-evasion problem where the pursuer has nonholonomic constraints. Since then, numerous works have studied pursuit-evasion games with dynamic constraints. In [9], the authors study the “Suicidal Pedestrian problem” inspired by Isaacs. Ruiz and Murrieta-Cid [10] analyze a scenario with a differential drive, faster evader against an omnidirectional but slower pursuer. In [11] Scott and Leonard derive and analyze optimal control strategies for speed, turning rate, lateral acceleration and movement direction-constrained evaders. In [12] the authors propose a single-agent reinforcement learning approach to the multi-pursuer with unicycle constraints scenario against a single, omnidirectional evader. Similarly, in [13] Yang et al. study a multi-pursuer, multi-evader version of pursuit-evasion problem with unicycle constraints and propose a MARL state processing method to improve the scalability.
@@
If the players can not observe each other at all times, for example, if the game is played in a complex environment under sensing limitations, pursuit-evasion game has been defined as a combination of two subcomponents: 1) locating the evader (search problem) and 2) capturing the evader. From the pursuer perspective, the search problem appears once sensor constraints are introduced to the pursuit-evasion problem. From the evader perspective, optimal strategies depend on the environment as well as the sensor constraints of the agents in the pursuit-evasion game. The problem of planning paths to reach a state with unobstructed view of the evader was first presented in [14]. In [1], it was shown that in a simply connected polygon of $n$ vertices, $\Theta(\log{n})$ pursuers are necessary and sufficient to detect an unpredictable evader. Numerous studies since then have formulated different variations of the pursuit-evasion problem, [15, 16, 17] with [15] studying the case with a pursuer with limited field of view and [16] with unreliable sensor data. In [2] Isler et al. propose a randomized algorithm to locate the evader with a single pursuer and extend this strategy to two pursuers with line-of-sight visibility to capture the evader in simply connected polygons. In later work, [18] Noori and Isler have shown that a single pursuer can locate and capture the evader in monotone polygons. In [19] Engin et al. have proposed a single-agent reinforcement learning method with a compact belief state representation for pursuit-evasion games with visibility constraints in simply-connected polygons. In [20] Bajcsy et al. have studied the pursuit-evasion problem for quadrupedal robots under real-world physical and visibility constraints and proposed a teacher-student network pair to deal with the issues caused by partial observability in the problem definition.
In this work, we investigate a game between players with complex dynamics and with sensing limitations.
@@ III-A Agent and Environment Models
$\dot{x_{1}}=x_{4}cos(x_{5})$ (1)
$\dot{x_{2}}=x_{4}sin(x_{5})$
$\dot{x_{3}}=f_{steer}(x_{3},u_{1})$
$\dot{x_{4}}=f_{acc}(x_{4},u_{2})$
$\dot{x_{5}}=\frac{x_{4}}{l_{f}+l_{r}}tan(x_{3})$
where $s_{x}$ and $s_{y}$ are the coordinates in meters in the global frame. $v$ is the longitudinal velocity in $m/s$, and $\Psi$ is the global yaw angle at the vehicle center in radians. Inputs are defined as $u_{1}$ representing the steering velocity $\dot{\delta}$ in $rad/s$ and $u_{2}$ representing the longitudinal acceleration $\dot{v}$ in $m/s^{2}$. System steering and acceleration input constraints are represented as $f_{steer}$ and $f_{acc}$, respectively. $l_{f}+l_{r}$ is the car model parameter representing the vehicle wheelbase with $l_{f}$ representing the distance of the front axle from the center of gravity and $l_{r}$ representing the same for the rear axle in meters. See Fig. 2a.
The evader is based on the point-mass state-space model [35] defined in Eq. (2), consisting of the following states $x_{1}=s_{x},x_{2}=s_{y},x_{3}=\dot{s_{x}},x_{4}=\dot{s_{y}}$ and uses the same global frame of reference notation with the inputs $u_{1}$ and $u_{1}$ representing the accelerations along the global x and y-axes $\ddot{s_{x}}$ and $\ddot{s_{y}}$ in $m/s^{2}$.
$\dot{x_{1}}=x_{3}$ (2)
$\dot{x_{2}}=x_{4}$
$\dot{x_{3}}=u_{1}$
$\dot{x_{4}}=u_{2}$
Both the pursuer and evader agents use the same sensor model represented by a wedge (slice of a circle) centered at each agent’s center of mass and have two parameters: field of view angle ($rad$) $V^{i}_{\alpha}$ and range ($m$) $V^{i}_{r}$ as shown in Fig. 2b, which may differ between the agents. An agent sees its opponent if and only if the opponent is inside the agent sensor footprint. The sensor footprint of an agent i is denoted by $V^{i}\subseteq\mathcal{E}$ with $i\in\{p,e\}$. Therefore, the pursuer $p$ sees the evader $e$ if $e\in V^{p}$ and vice versa. The sensor footprint of an agent is also presented in Fig. 2b.
@@ III-B Game Formulation
Within the scope of this paper, the pursuit evasion game is defined as a POSG. However, it must be noted that without added stochastic noise, presented agent dynamics are fully deterministic as described in Eq. 1 and Eq. 2. The formal definition of a POSG is shown in 1 as defined by Terry et al. [36] inspired by Shapley [21].
Definition 1
A Partially Observable Stochastic Game (POSG) is a tuple $\langle\mathcal{S},s_{0},N,A^{i},P,R^{i},\Omega^{i},O^{i}\rangle$, where:
  • $\mathcal{S}$ denoting the set of possible states.
  • $s_{0}$ denoting the initial state.
  • $N$ is the number of agents. The set of agents is denoted as $[N]$
  • $\mathcal{A}^{i}$ is the set of possible actions for agent $i$ with $i\in[N]$.
  • $P:\mathcal{S}\times\prod_{i}\mathcal{A}^{i}\times\mathcal{S}$$\rightarrow[0,1]$ is the transition function. It has the property that for all $s\in\mathcal{S}$, for all $(a^{1},a^{2},\ldots,a^{N})\in\prod_{i}\mathcal{A}^{i}$,$\sum_{s^{\prime}\in\mathcal{S}}P(s,a^{1},a^{2},\ldots,a^{N},s^{\prime})=1.$
  • $R^{i}:\mathcal{S}\times\prod_{i}\mathcal{A}^{i}\times\mathcal{S}\rightarrow\mathbb{R}$ is the reward function for agent $i$.
  • $\Omega^{i}$ is the set of possible observations for agent i.
  • $O^{i}:\mathcal{A}^{i}\times\mathcal{S}\times\Omega^{i}\rightarrow[0,1]$ is the observation function. It has property that $\sum_{\omega\in\Omega^{i}}O^{i}(a,s,\omega)=1$ for all $a\in\mathcal{A}^{i}$ and $s\in\mathcal{S}$
@@ IV Method
The objective of each agent is to maximize its total expected reward $G^{i}=\sum_{t=0}^{\mathcal{T}}{\gamma_{t}r^{i}_{t}}$ over time horizon $\mathcal{T}$ where $r^{i}_{t}$ is provided by the environment at timestep $t$ and $\gamma\in[0,1]$ is the discount factor. Per-agent actions are sampled from the agent policy networks $a^{i}_{t}\sim\pi^{i}_{\theta}(s^{i}_{t})$.
@@ IV-A Multi-Agent Deep Reinforcement Learning
We use the MADDPG algorithm [5] which follows the Centralized Critic, Decentralized Execution (CTDE) principle to train the agents by making them play against each other from scratch.
@@ IV-B Space Representations
The hidden full state of the pursuit-evasion game between a kinematic car-like pursuer and a point-mass evader is referred as a 9-vector $s\coloneqq[x^{p}_{1},x^{p}_{2},x^{p}_{3},x^{p}_{4},x^{p}_{5},x^{e}_{1},x^{e}_{2},x^{e}_{3},x^{e}_{4}]^{\mathsf{T}}$. Without loss of generality, the state vector comprises of state variables presented in Section III-A. Time indices are dropped for brevity. Each agent in the pursuit-evasion game creates its own observation vector $o^{i}$ through partial observations from the environment represented as 11-vectors of floating point numbers by appending a binary observation flag (represented with floating point number) indicating the visibility of the opponent and a time index. If an agent cannot observe its opponent at a timestep, corresponding values in $o^{i}$ are zeroed and the binary flag is set to negative. Observation vectors are normalized to contain values in interval $[-1,1]$. Input/action representations follow the same convention described in III-A and are represented as 2-vectors normalized to the interval $[-1,1]$ by taking acceleration, velocity, steering angle velocity and steering angle limits. Both the state and the action spaces are continuous.
IV-C Reward Structure

### v2

# Pursuit-Evasion for Car-like Robots with Sensor Constraints

@@ Abstract
We study a pursuit-evasion game between two players with car-like dynamics and sensing limitations by formalizing it as a partially observable stochastic zero-sum game. The partial observability caused by the sensing constraints is particularly challenging. As an example, in a situation where the agents have no visibility of each other, they would need to extract information from their sensor coverage history to reason about potential locations of their opponents. However, keeping historical information greatly increases the size of the state space. To mitigate the challenges encountered with such partially observable problems, we develop a new learning-based method that encodes historical information to a belief state and uses it to generate agent actions. Through experiments we show that the learned strategies improve over existing multi-agent RL baselines by up to 16% in terms of capture rate for the pursuer. Additionally, we present experimental results showing that learned belief states are strong state estimators for extending existing game theory solvers and demonstrate our method’s competitiveness for problems where existing fully observable game theory solvers are computationally feasible. Finally, we deploy the learned policies on physical robots for a game between the F1TENTH and JetRacer platforms moving as fast as 2 m/s in indoor environments, showing that they can be executed on real-robots.
@@ I Introduction
As an example, in the classical homicidal chauffeur game [3], the pursuer is a car-like vehicle whose goal is to hit a pedestrian who is not subject to any turning constraints. In this game, the pursuer has three state variables (position and orientation) and the evader has two (only position). One might view the solution of this game as a subspace (manifold) of this five dimensional joint state space. The tangent space of this solution manifold corresponds to the optimal control actions of the players. In some cases, the differential equations for the evolution of the game when the players play optimally can be derived in closed form. The strategies for a given instance are usually obtained by numerical integration. Alternatively, the state space can be discretized and the solution can be obtained using dynamic programming.
Figure 1: Real-world experiments: Execution of the learned pursuit-evasion policies on two robotic cars with different wheelbase, velocity and steering angle constraints.
However, such methods become quickly intractable as the dimensionality of the game increases. For example, in his book, Isaacs presents only a partial solution for the “game of two cars” where both players have car-like dynamics. The dimensionality issue becomes more severe when sensing is involved; in the standard formulations, it is assumed that the players can observe each others’ state fully at each time step. This may not be always true due to observability limitations. The resulting “partial observability games” have even higher (sometimes infinite) dimensionality as the players now need to reason about all feasible states that are consistent with their observation histories.
Reinforcement Learning (RL) provides an alternative approach to construct the solution manifold by learning it through game play. In this paper, we extend recently developed multi-agent reinforcement learning (MARL) techniques by combining Mixture Density Networks [4] and bidirectional Long Short-term Memory (LSTM) Networks [5, 6, 7] to build a novel belief network as shown in Fig. 2, named BiMDN. We demonstrate that our proposed method can solve a challenging game where the pursuer is a car with five state variables including velocities better than existing MARL algorithms by encoding agent observation histories to belief states. We investigate two versions of the evader. It is either a car with the same kinematic model as the pursuer or a point-mass. Hence, the joint-state dimensionality can be as high as 10. The players are subject to field-of-view sensing restrictions for both the viewing angle and viewing distance (Section III). We cast the game as a competitive zero-sum Partially Observable Stochastic Game (POSG) and use the MARL algorithm multi-agent deep deterministic policy gradient (MADDPG) [8] with BiMDN to simultaneously train pursuer and evader policies. Our problem formulation and proposed approach do not require any pretrained or designed expert evader policy and generalize over varying agent constraints.
@@
  • We introduce a novel belief network named BiMDN for MARL to solve a pursuit-evasion game with complex dynamics and sensing constraints. We show that our method is able to train stronger pursuer policies without the need for expert policies even as compared to vanilla MARL algorithms.
  • In simulation experiments we compare our pursuer and evader policies to differential game solver baselines from the reachability literature and demonstrate competitive performance for simpler problems where existing solvers are computationally feasible.
  • We deploy the learned policies on physical robots for a game between the F1TENTH[9] and JetRacer [10] platforms moving at high-speed and showcase a successful direct policy transfer to real world.
Overall, our results show the potential of recurrent mixture models coupled with multi-agent reinforcement learning in navigating complex pursuit-evasion scenarios within dynamic environments and demonstrate that simulation-trained policies are directly transferable to real autonomous agents.
Figure 2: Our method: Agents receive partial observations $o^{i}_{t}$ at each timestep and store observation memories. For each agent, our BiMDN models encode individual agent observation histories to belief states $b_{t}^{i}$, whose uncertainty can be explicitly quantified. Trained actor networks $\mu_{\theta}^{i}$ use combined information from the belief states and environment partial observations to compute actions $a^{i}_{t}$.
@@ II-A Pursuit-Evasion
The mathematical origins of the pursuit-evasion problem can be traced back to 1930’s as the Lion and Man problem. In the original formulation, a lion and a human are in a circular area where the lion tries to catch an escaping human as quickly as possible. Various variants of this game have been studied in the robotics literature [11]. In his seminal work [3] Isaacs presented a differential game based formulation to the pursuit-evasion problem which has served as the basis of game-theoretic approaches.
@@
In control-theoretic formulations, both of the agents in a pursuit-evasion problem have constraints on maximum velocities. One prominent example is the Homicidal chauffeur problem formulated by Isaacs [3], which is a special case of the pursuit-evasion problem where the pursuer has nonholonomic constraints. Since then, numerous works have studied pursuit-evasion games with dynamic constraints[12, 13, 14]. In [15] the authors propose a single-agent reinforcement learning approach to the multi-pursuer with unicycle constraints scenario against a single, omnidirectional evader. Similarly, in [16] Yang et al. study a multi-pursuer, multi-evader version of pursuit-evasion problem with unicycle constraints.
Reachability literature presents solutions to various pursuit-evasion problems. ToolboxLS [17] demonstrates an Hamilton-Jacobi (HJ) solver by formulating pursuit-evasion as a differential game. HelperOC [18] and OptimizedDP [19] establish further improvements in performance and handling a higher number of dimensions for HJ solvers. Similarly, game-theoretic studies also treat pursuit-evasion games as general-sum games, provide various solvers [20, 21, 22, 23, 24, 25] searching for Nash equilibria.
@@
If the players can not observe each other at all times, pursuit-evasion game has been defined as a combination of two subcomponents: 1) locating the evader (search problem) and 2) capturing the evader. From the pursuer perspective, the search problem appears once sensor constraints are introduced to the pursuit-evasion problem. From the evader perspective, optimal strategies depend on the environment as well as the sensor constraints of the agents in the pursuit-evasion game. The problem of planning paths to reach a state with unobstructed view of the evader was first presented in [26]. In [1], it was shown that in a simply connected polygon of $n$ vertices, $\Theta(\log{n})$ pursuers are necessary and sufficient to detect an unpredictable evader. Numerous studies since then have formulated different variations of the pursuit-evasion problem, [27, 28, 29] with [27] studying the case with a pursuer with limited field of view and [28] with unreliable sensor data. In [2] Isler et al. propose a randomized algorithm to locate the evader with a single pursuer and extend this strategy to two pursuers with line-of-sight visibility to capture the evader in simply connected polygons. In later work, [30] Noori and Isler have shown that a single pursuer can locate and capture the evader in monotone polygons. In [31] Engin et al. have proposed a single-agent reinforcement learning method with a compact belief representation for pursuit-evasion games with visibility constraints in simply-connected polygons. In [32] Zhang et al. propose a MARL-based approach for n-pursuer pursuit-evasion game between drones in urban maps, incorporating an LSTM network for future state predictions in simulated experiments. In [33] Bajcsy et al. have studied the pursuit-evasion problem for quadrupedal robots under real-world physical and visibility constraints.
In this work, we differentiate ourselves from existing literature by studying a game between players with complex dynamics and sensing limitations; assuming no prior knowledge of the dynamics or the sensor model, which are typical limitations of the methods proposed in the reachability and game theory solver literature. Additionally, unlike some of the prior works, we do not train against an expert policy to showcase competitive performance in simulation as well as real robot experiments.
@@ II-B Dynamic Games & MARL
II-C Belief States
A significant number of problems in robotics domain inherently present some form of partial observability which breaks Markovian assumption made by standard RL literature. There is a body of research proposing the use of belief states [48, 49, 50], which are observations conditioned on the history of interactions to preserve Markovian properties. Contemporary partially observable RL literature presents various recurrent models for belief modeling [51, 52, 53, 54], aiming to capture spatiotemporal features relevant to the downstream task. With BiMDN, we additionally focus on modeling the spatial uncertainty presented by the pursuit-evasion problem to learn stronger policies in multi-agent problems.
@@ III-A Agent and Environment Models
$\dot{x_{1}}=x_{4}\cos(x_{5})$ $\dot{x_{2}}=x_{4}\sin(x_{5})$ (1)
$\dot{x_{3}}=f_{\mathrm{steer}}(x_{3},u_{1})$ $\dot{x_{4}}=f_{\mathrm{acc}}(x_{4},u_{2})$
$\dot{x_{5}}=\frac{x_{4}}{l_{f}+l_{r}}\tan(x_{3})$
where $s_{x}$ and $s_{y}$ are the coordinates in meters in the global frame. $v$ is the longitudinal velocity in $m/s$, and $\Psi$ is the global yaw angle at the vehicle center in radians. Inputs are defined as $u_{1}$ representing the steering velocity $\dot{\delta}$ in $rad/s$ and $u_{2}$ representing the longitudinal acceleration $\dot{v}$ in $m/s^{2}$. System steering and acceleration input constraints are represented as $f_{steer}$ and $f_{acc}$, respectively. $l_{f}+l_{r}$ is the car model parameter representing the vehicle wheelbase with $l_{f}$ representing the distance of the front axle from the center of gravity and $l_{r}$ representing the same for the rear axle in meters with zero slip angle assumption. See Fig. 3a.
The evader is based on the point-mass model [55] defined in Eq. (2), consisting of the following states $x_{1}=s_{x},x_{2}=s_{y},x_{3}=\dot{s_{x}},x_{4}=\dot{s_{y}}$ and uses the same global frame of reference notation with the inputs $u_{1}$ and $u_{2}$ representing the accelerations along the global x and y-axes $\ddot{s_{x}}$ and $\ddot{s_{y}}$ in $m/s^{2}$.
$\dot{x_{1}}=x_{3},\quad\dot{x_{2}}=x_{4},\quad\dot{x_{3}}=u_{1},\quad\dot{x_{4}}=u_{2}$ (2)
@@
Pursuer agent uses the noise-free sensor model represented by a wedge (slice of a circle) centered at agent center of mass and have two parameters: field of view angle ($rad$) $V^{i}_{\alpha}$ and range ($m$) $V^{i}_{r}$ as shown in Fig. 3b. An agent sees its opponent if and only if the opponent is inside the agent sensor footprint. The sensor footprint of an agent i is denoted by $V^{i}\subseteq\mathcal{E}$ with $i\in\{p,e\}$. Therefore, the pursuer $p$ sees the evader $e$ if $e\in V^{p}$ and vice versa. The sensor footprint of an agent is presented in Fig. 3b and the visibility condition is also defined as follows with the help of a binary visibility variable $\mathbb{I}$:
$\mathbb{I}^{i}=\begin{cases}+1,&\text{if }d\bigl{(}(x_{1}^{i},x_{2}^{i}),(x_{1}^{j},x_{2}^{j})\bigr{)}\leq V^{i}_{r}\text{ and }\\ &\quad\frac{-V^{i}_{\alpha}}{2}\leq\operatorname{atan2}(x_{2}^{j}-x_{2}^{i},x_{1}^{j}-x_{1}^{i})-x^{i}_{5}\leq\frac{V^{i}_{\alpha}}{2},\\ -1,&\text{otherwise}.\end{cases}$
where $i$ and $j$ denote an agent and its opponent, respectively. Angular differences are wrapped to the interval $[-\pi,\pi)$. Measurement function is defined formally as follows :
$\mathbb{M}^{i}=\begin{cases}\mathbf{x}^{j},&\text{if }\mathbb{I}^{i}=+1,\\ \mathbf{0}\in\mathbb{R}^{n},&\text{otherwise}.\end{cases}$ (3)
where the function returns a vector of opponent states or zeros depending on the observability. Since point-mass agents don’t have a heading, $V^{i}_{\alpha}$ is always $2\pi\ rad$ for point-mass agents, making the angular constraint negligible and forming a disk with the footprint.
@@ III-B Game Formulation
Within the scope of this paper, the pursuit evasion game is defined as a POSG. However, it must be noted that without added stochastic noise, presented agent dynamics are fully deterministic as described in Eq. 1 and Eq. 2.
@@ IV Method
The objective of each agent is to maximize its total expected reward $G^{i}=\sum_{t=0}^{\mathcal{T}}{\gamma_{t}r^{i}_{t}}$ over time horizon $\mathcal{T}$ where $r^{i}_{t}$ is provided by the environment at timestep $t$ and $\gamma\in[0,1]$ is the discount factor. Per-agent actions are sampled from the agent policy networks $a^{i}_{t}\sim\mu^{i}_{\theta}(s^{i}_{t})$.
@@ IV-A Multi-Agent Deep Reinforcement Learning
We use the MADDPG algorithm [8] which follows the Centralized Critic, Decentralized Execution (CTDE) principle to train the agents by making them play against each other from scratch with access to privileged information on the opponents during training. This is different from self-play, where agents are completely independent during the training stage without access to privileged information. Our method is also model-free, i.e. agents are not provided with information on the dynamic equations that govern the interactions or learn such equations. Instead, both agents train corresponding value and policy networks through interactions.
@@ IV-B Space Representations
The hidden full state of the pursuit-evasion game between a kinematic car-like pursuer and a point-mass evader is referred as a 9-vector $s\coloneqq[x^{p}_{1},x^{p}_{2},x^{p}_{3},x^{p}_{4},x^{p}_{5},x^{e}_{1},x^{e}_{2},x^{e}_{3},x^{e}_{4}]^{\mathsf{T}}$. Without loss of generality, the state vector comprises of state variables presented in Section III-A. Time indices are dropped for brevity. Each agent in the pursuit-evasion game creates its own observation vector $o^{i}$ through partial observations from the environment represented as 11-vectors of floating point numbers by appending a binary observation flag $\mathbb{I}^{i}$ indicating the visibility of the opponent and a time index. If an agent cannot observe its opponent at a timestep, corresponding values in $o^{i}$ are zeroed and the binary flag $\mathbb{I}^{i}$ is set to $-1$. Observation vectors are normalized to contain values in interval $[-1,1]$. Input/action representations follow the same convention described in III-A and are represented as 2-vectors normalized to the interval $[-1,1]$ by taking acceleration, velocity, steering angle velocity and steering angle limits. Both the state and the action spaces are continuous.
IV-C Belief Network
We extend each agent’s observation vector with information derived from belief states $b^{i}_{t}$. BiMDN is a sequential model of bidirectional LSTM layers followed by fully connected layers and output heads for predicting a Mixture of Gaussians parameterized by the means $M^{i}_{t}$, standard deviations $\Sigma^{i}_{t}$ and cluster weights $\Pi^{i}_{t}$ using past and current observations. Predictions are made over 2 variables which represent the agent beliefs along the $x$ and $y$ axes. Prediction clusters are axis-aligned ellipsoids which is equivalent to a diagonal covariance matrix as we have experienced no significant improvement in performance by further parametrization of the covariance matrix. Similar to the CTDE paradigm (or asymmetric RL), training of BiMDN uses privileged information as ground truth with Negative Log Likelihood loss derived from $M^{i}_{t}$, $\Sigma^{i}_{t}$ and $\Pi^{i}_{t}$.
BiMDN learns to encode relevant aspects from a time series input of stacked historical observation vectors to a compact belief state. During training, unlike commonly used neural network architectures like LSTM networks, our architecture explicity quantifies uncertainty about its belief and does not output point estimates. This is a useful property in pursuit-evasion games where an average point estimate would be incapable of modeling the potential regions the opponent may be at. To indicate a higher uncertainty, the clusters expand in size proportional to their standard deviations and get dispersed across the environment. Further discussion including network parameters is provided in Section V-A.
IV-D Reward Structure